OpenStudy (anonymous):

Solve the equation: x2 - 12x = -36

OpenStudy (mimi_x3):

\[x^2-12x+36\] \(a=1,b=-12,c=36\) then use the quadratic formula \[x=\frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2a} \]

OpenStudy (mimi_x3):

\(x^2-12x+36=0\)*

OpenStudy (anonymous):

it's factorable -6,-6 =P

OpenStudy (apoorvk):

Why can;t we just factor this by the way?? :o \(x^2 - 6x - 6x +36\) \(=x(x-6)-6(x-6)\) \(=(x-6)^2\)

OpenStudy (mimi_x3):

yeah lol. \(x^2-12x=-36\) => \(x^2-12x+36\) \(a+b= -6+-6\) \(a*b=6*6\) you can justuse the quadratic formula anyway; its faster.